OpenStudy (anonymous):

What is the square of 3√b?

OpenStudy (anonymous):

cancel the square root and square the 3

OpenStudy (anonymous):

(3sqrt(b))^2=3^2*sqrt(b)^2=9*b

OpenStudy (anonymous):